OpenStudy (anonymous):

which ordered pair is a solution of the equation y=-9x+4 fan and gold reward

OpenStudy (anonymous):

Solutions?

OpenStudy (anonymous):

a (10,-86) b (-4,-58) c (6,-41) d (-6, 57)

OpenStudy (anonymous):

-9(10)+4=-86 -90+4=-86 -86=-86 (10,-86)

OpenStudy (anonymous):

thank you so much. can i ask u a few more ?s

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

solve the following equation. -3=z+4 z=4 z=-4 z=-7 z=7

OpenStudy (anonymous):

w-2=-3 w=3 - 2 w= -5 w=-1 w=6

OpenStudy (anonymous):

-4+-3=z -7=z

OpenStudy (anonymous):

w-2=-3 W-2+2=-3+2 w=-1

OpenStudy (anonymous):

wow you are amazing 0_/

OpenStudy (anonymous):

To solve, get the variable by itself by moving the number to the other side.

OpenStudy (anonymous):

thank you